[Brahms-l] Shift report 20050323 16:00-24:00

From: <apache@brahms-web.brahms.bnl.gov>
Date: Wed Mar 23 2005 - 00:08:25 EST
Shifter leader was sjs
Shifters were 


Started shift preparing for injection voltage.  FV fixed timing on triggers.  Because of rapid falloff of beam intensity, changing runs every 1/2 hour.  Also, we are not lowing voltages on beam dump.
FS @ 4deg B861; MRS @ 40 deg  B700
Run    Trig2
13925   2.8K
13926   4.4K
13927  10.2K

FS @ 8deg B861;  MRS @ 40 deg B700 (FS fields should have been lowered)
Run     Trig2
13928    4.7K
13929    2.8K
13930    1.6K

FS @8deg B430;  MRS @ 40 deg B700
Run     Trig2
13931     0.9K
13934     7.3K
13935     0.4K
13936     0.4K
13937     3.1K
13938     6.6K
13939     3.8K
13940     0.4K
13941     6.6K
13942  running

pedestal runs: 13932 and 13933

No new detector problems noted.

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Wed Mar 23 00:12:05 2005

This archive was generated by hypermail 2.1.8 : Wed Mar 23 2005 - 00:12:14 EST